Find the coordinates of the x and y intercepts
x-10=4y

Answer by nerdybill(7384)

To find the x-intercepts, set y=0 and solve for x:
x-10=4y
x-10=4(0)
x-10=0
x = 10
x-intercept is at (10,0)
.
To find the y-intercepts, set x=0 and solve for y:
x-10=4y
0-10=4y
-10/4=y
-5/2 = y
y-intercept is at (0, -5/2)
.